# Equilibrium-distribution-function based mesoscopic finite-difference methods for partial differential equations: Modeling and Analysis

### v1

@@ III.2 The MesoFD and MMFD schemes with a three-level explicit structure for PDE
For simplicity, take $k\in\{-1,0,1\}$, and let $a_{k0}=a_{k},a_{k,-1}=\hat{a}_{k},b_{k0}=b_{k},b_{k,-1}=\hat{b}_{k}$, then Eqs. (46) and (47) become
@@
$\hat{a}_{1}^{+}=\hat{a}_{1}+\hat{a}_{-1},\hat{a}_{1}^{-}=\hat{a}_{1}-\hat{a}_{-1},\hat{b}_{1}^{+}=\hat{b}_{1}+\hat{b}_{-1},\hat{b}_{1}^{-}=\hat{b}_{1}-\hat{b}_{-1}.$ (51)
According to Remark 1, for Eqs. (48), (49) and (III.2), Eq. (III.1) has the following form
@@
where we can obtain the relation $A_{20}=-A_{10}/2-1$.
@@
(1) SRT LB model with $\tau=1$. Set $a_{0}=a_{1}=0,a_{-1}=1;\hat{a}_{0}=\hat{a}_{1}=\hat{a}_{-1}=0$. $b_{0}=b_{1}=\hat{b}_{0}=\hat{b}_{1}=0,b_{-1}=3/2,\hat{b}_{-1}=-1/2$.
(2) TRT LB model with $s^{-}+s^{+}=2$[9, 10, 22]. Set $a_{0}=0,a_{1}=1-s^{-},a_{-1}=1;\hat{a}_{0}=-(1-s^{-}),\hat{a}_{1}=\hat{a}_{-1}=0$. $b_{0}=\hat{b}_{0}=-(1-s^{-})/2,b_{1}=\hat{b}_{1}=0,b_{-1}=3/2,\hat{b}_{-1}=-1/2$.
(3) One-stage simplified LB model without source term[25, 26]. Set $a_{0}=2(1-\tau),a_{1}=\tau-1,a_{-1}=\tau;\hat{a}_{0}=\hat{a}_{1}=\hat{a}_{-1}=0$.
(4) A new FD scheme for the wave equation without source term (Eq. (45)). Set $a_{0}=2(1-\gamma),a_{1}=a_{-1}=\gamma;\hat{a}_{0}=-1,\hat{a}_{1}=\hat{a}_{-1}=0$, where $\gamma=\alpha/(\beta c_{s}^{2})$.
In general, to derive a FD scheme one can determine the parameters with Eq. (52). We now show that Eq. (52a) is a solvable linear system. In fact, let $X=(a_{1},a_{-1},\hat{a}_{0},\hat{a}_{-1},\hat{a}_{-1})^{T}$, $b=(1-a_{0},1+A_{10},A_{11},A_{21},A_{22})^{T}$, then Eq.(52a) can be transformed into a solvable linear system $AX=b$ with the invertible coefficient matrix $A$,
@@
III.3 The stability of mesoscopic numerical method based macroscopic finite difference scheme for linear CDE
In this part, we consider the stability of MMFD scheme (4) for the linear CDE (LCDE) without the source term. The LCDE has the following form,
@@
where $\mathbf{u}$ is a constant vector, and $\mathbf{D}_{0}$ is a constant matrix (tensor). When $\mathbf{D}_{0}=\mathbf{0}$ the LCDE above becomes a CE. The MesoFD scheme for Eq. (56) can be simplified as
@@
where the $\omega_{j}$ is the weight coefficient (see Appendix A for details), and $\mathbf{C}=c_{s}^{2}(\mathbf{\beta D}_{0}-\mathbf{I})+\bar{\lambda}\mathbf{uu},~{}\mathbf{Q}_{j}=\mathbf{c}_{j}\mathbf{c}_{j}-c_{s}^{2}\mathbf{I},~{}c_{s}^{2}=d_{0\alpha}c_{\alpha}^{2},~{}c_{\alpha}=\Delta x_{\alpha}/\Delta t$, and $\bar{\lambda}=0$ or 1, then summing Eq. (58), we obtain
@@
$\hat{\phi}(t)=\sum_{q}A_{q}\hat{\phi}(t+q\Delta t),A_{q}=\sum_{k}a_{kq}\sum_{j\geq 0}\big{[}\bar{u}_{j}^{+}\cos(k\xi_{j})+\textbf{i}\bar{u}_{j}^{-}\sin(k\xi_{j})\big{]},~{}q\leq 0,$ (62)
or equivalently
$\hat{\phi}(t)=\sum_{q<0}\frac{A_{q}}{1-A_{0}}\hat{\phi}(t+q\Delta t),~{}A_{q}=\sum_{k}a_{kq}\sum_{j\geq 0}\big{[}\bar{u}_{j}^{+}\cos(k\xi_{j})+\textbf{i}\bar{u}_{j}^{-}\sin(k\xi_{j})\big{]},~{}q\leq 0.$ (63)
For a two-level scheme with $q\in\{0,-1\}$, Eq. (63) is simplified as
$\hat{\phi}(t)=\frac{A_{-1}}{1-A_{0}}\hat{\phi}(t-\Delta t),A_{q}=\sum_{k}a_{kq}\sum_{j\geq 0}\big{[}\bar{u}_{j}^{+}\cos(k\xi_{j})+\textbf{i}\bar{u}_{j}^{-}\sin(k\xi_{j})\big{]},~{}q=0,-1,$ (64)
with the stability condition
@@
where $a_{k}=a_{k,-1},a_{-k}=a_{-k,-1},k=1,...,m$, and $m=m_{-1}$.
Note that Eq. (III.3) has the same form as Eq. (77) in Ref. [40]. Based on Eq. (III.3) the stability theorem in Ref. [40] can be extended as follows.
Theorem 1 The two-level explicit scheme is stable ($|\lambda|\leq 1$) if and only if the following conditions holds
@@
$\lambda=\frac{1+\theta(A+iB)}{1-(1-\theta)(A+iB)},$ (71)
@@
Theorem 2 The two level $\theta$-scheme (Eq. (II.2)) without the source term is (1) unconditional stable ($|\lambda|\leq 1$) for $20\leq\theta\leq 1/$ 2, (2) conditional stable for $\theta>1/2$ if Eq. (67a) holds.
Taking $\mathbf{D}_{0}=\mathbf{I},\beta=1,\bar{\lambda}=0$ in Eq. (59), the linear EDF $f_{j}^{eq}$ with $\bar{u}_{j}^{+}=\omega_{j}>0$ and $\bar{u}_{j}^{-}=\omega_{j}\frac{\mathbf{c}_{j}\cdot\mathbf{u}}{c_{s}^{2}}$ is obtained, and
@@
Further the equivalent form of Eq. (57) can be derived from Eq. (16) as
@@
For the two-level explicit scheme, Eq. (III.3) is simplified as
@@
where Eq. (III.1), Eq. (40) and Eq. (75b) are used. Note that Eq. (79b) is equivalent to Eq. (75b) or (76). When $m=1$, Eq. (79a) is simplified as
@@
and the stability condition Eq. (75) becomes
@@
and the stability condition Eq. (81) is equivalent to
@@
One can find that the condition in Eq. (86b) is much stronger than that in Eq. (87) for large $d$.
@@ III.4 The derivation of the moment equations for NSEs using DTE method
where $\mathbf{M}_{k}$ and $\mathbf{M}_{kF}$ ($k\geq 0$) are the $k$-th moments of $f_{j},f_{j}^{eq}$ and $F_{j}$, respectively. Additionally, Eq. (88a) gives the following moments of nonequilibrium,
@@
Similarly, using DTE method we can derive the moment equations based on Eq.(1) or Eq. (5). Before doing that, we first use the double distribution based scheme Eq. (5) in the form of moments:
@@
in which two different set of parameters, $\{a_{kq},b_{kq}\}$ and $\{\bar{a}_{kq},\bar{b}_{kq}\}$ are used.
Using Taylor expansion to Eq.(90) we have
@@
Taking $\bar{A}_{00}=A_{00}=1$ in Eq. (91), we obtain a general moment equation for NSEs
@@
Note that Eq. (93a) is same as Eq. (30). In what follows, we only discuss Eq.(93b) which gives for $N=1,2$ that
@@
the moment equation in the leading order of $O(\Delta t^{2})$ can be derived from Eq. (96b),
@@
Further according to a given NSEs, and choosing the moments and parameters in Eqs. (39) and (99) properly, we can correctly recover the NSEs from Eqs. (39) (or (100)) and (99). For example, taking
@@
in Eqs. (100) and (99) leads to the general NSEs
@@
where the auxiliary source term $\mathbf{M}_{30}$ contains the additional term caused by the anisotropy of lattice tensor in a given lattice model as mentioned in Ref. [17]. If an isotropic lattice is used, $\mathbf{M}_{30}=\mathbf{0}$ can be taken.
@@
therefore, the incompressible NSEs are recovered from Eqs. (103) and (102(b)) in the weak compressible limit, and $\mathbf{M}_{2F}$ can be simplified [17].

### v2

@@ III.2 The MesoFD and MMFD schemes with a three-level explicit structure for PDE
For simplicity, if we take $k\in\{-1,0,1\}$, and let $a_{k0}=a_{k},a_{k,-1}=\hat{a}_{k},b_{k0}=b_{k},b_{k,-1}=\hat{b}_{k}$, then Eqs. (48) and (49) become
@@
$\hat{a}_{1}^{+}=\hat{a}_{1}+\hat{a}_{-1},\hat{a}_{1}^{-}=\hat{a}_{1}-\hat{a}_{-1},\hat{b}_{1}^{+}=\hat{b}_{1}+\hat{b}_{-1},\hat{b}_{1}^{-}=\hat{b}_{1}-\hat{b}_{-1}.$
According to Remark 1, for Eqs. (50), (51) and (III.2), Eq. (III.1) has the following form,
@@
from which we can obtain the relation $A_{20}=-A_{10}/2-1$.
@@
(1) SRT LB model with $\tau=1$: $a_{0}=a_{1}=0,a_{-1}=1;\hat{a}_{0}=\hat{a}_{1}=\hat{a}_{-1}=0$; $b_{0}=b_{1}=\hat{b}_{0}=\hat{b}_{1}=0,b_{-1}=3/2,\hat{b}_{-1}=-1/2$.
(2) TRT LB model with $s^{-}+s^{+}=2$ [9, 10, 22]: $a_{0}=0,a_{1}=1-s^{-},a_{-1}=1;\hat{a}_{0}=-(1-s^{-}),\hat{a}_{1}=\hat{a}_{-1}=0$; $b_{0}=\hat{b}_{0}=-(1-s^{-})/2,b_{1}=\hat{b}_{1}=0,b_{-1}=3/2,\hat{b}_{-1}=-1/2$.
(3) One-stage simplified LB model without source term [25, 26]: $a_{0}=2(1-\tau),a_{1}=\tau-1,a_{-1}=\tau;\hat{a}_{0}=\hat{a}_{1}=\hat{a}_{-1}=0$.
(4) A new FD scheme for the wave equation without source term (Eq. (47)): $a_{0}=2(1-\gamma),a_{1}=a_{-1}=\gamma;\hat{a}_{0}=-1,\hat{a}_{1}=\hat{a}_{-1}=0$, where $\gamma=\alpha/(\beta c_{s}^{2})$.
In general, to derive a FD scheme, one can determine the parameters with Eq. (54). We now show that Eq. (54a) is a solvable linear system. In fact, let $X=(a_{1},a_{-1},\hat{a}_{0},\hat{a}_{-1},\hat{a}_{-1})^{T}$, $b=(1-a_{0},1+A_{10},A_{11},A_{21},A_{22})^{T}$, then Eq.(54a) can be transformed into a solvable linear system $AX=b$ with the invertible coefficient matrix $A$,
@@
III.3 The stability of mesoscopic numerical method based macroscopic finite difference scheme for linear CDE and WE
In this part, we consider the stability of MMFD scheme (4) for the linear CDE (LCDE) and WE (LWE) without the source term, respectively. The LCDE and LWE have the following forms,
@@
$\partial_{t}^{2}\phi=\nabla\cdot\big{[}\alpha\nabla\cdot\mathbf{D}_{0}\phi\big{]},$ (59)
where $\mathbf{u}$ is a constant vector, and $\mathbf{D}_{0}$ is a constant matrix (tensor). When $\mathbf{D}_{0}=\mathbf{0}$, the LCDE will become a CE. The MesoFD scheme for Eq. (58) and Eq. (59) can be simplified as
@@
where $\omega_{j}$ is the weight coefficient (see Appendix A for details), $\mathbf{C}=c_{s}^{2}(\mathbf{\beta D}_{0}-\mathbf{I})+\bar{\lambda}\mathbf{uu},~{}\mathbf{Q}_{j}=\mathbf{c}_{j}\mathbf{c}_{j}-c_{s}^{2}\mathbf{I},~{}c_{s}^{2}=d_{0\alpha}c_{\alpha}^{2},~{}c_{\alpha}=\Delta x_{\alpha}/\Delta t$, and $\bar{\lambda}=0$ or 1. For the WE (59), we can set $\mathbf{u}=0$ in Eq. (62). Then summing Eq. (61), we obtain
@@
$\hat{\phi}(t)=\sum_{q=-n}^{0}A_{q}\hat{\phi}(t+q\Delta t),$ (65a)
$A_{q}$ $=$ $\sum_{k}a_{kq}\sum_{j\geq 0}\big{[}\bar{u}_{j}^{+}\cos(k\xi_{j})+\textbf{i}\bar{u}_{j}^{-}\sin(k\xi_{j})\big{]}$ (65b)
$=$ $\sum_{k}a_{kq}+\sum_{k\geq 1}\sum_{j\geq 0}\big{[}a_{kq}^{+}\bar{u}_{j}^{+}(\cos(k\xi_{j})-1)+\textbf{i}a_{kq}^{-}\bar{u}_{j}^{-}\sin(k\xi_{j})\big{]},~{}q\leq 0,$
where $\sum_{j}\bar{u}_{j}^{+}=1$ has been used, $a_{kq}^{+}anda_{kq}^{-}$ are defined by Eq. (14). Actually, Eq. (65a) is equivalent to
$\hat{\phi}(t)=\sum_{q=-n}^{-1}\frac{A_{q}}{1-A_{0}}\hat{\phi}(t+q\Delta t).$ (66)
The characteristic polynomial of Eq. (66) is of degree $n$, and has the following form
$p(\lambda)=\lambda^{n}-\sum_{k=0}^{n-1}p_{k-n}\lambda^{k},~{}p_{q}=\frac{A_{q}}{1-A_{0}},~{}q=-n,~{}...,~{}-1,$ (67)
the MMFD scheme is stable if the modulus of any root $\lambda$ of this polynomial is no greater than 1. Usually, when $n$ is large, one cannot easily obtain the stability condition. For this reason, here we consider some special cases.
For a two-level scheme with $q\in\{0,-1\}$, Eq. (66) can be simplified as
$\hat{\phi}(t)=\frac{A_{-1}}{1-A_{0}}\hat{\phi}(t-\Delta t),$ (68)
and the stability condition is given by
@@
where $\sum_{k}a_{k,-1}=1$ is used, $a_{k}=a_{k,-1},a_{-k}=a_{-k,-1},k=1,...,m$, and $m=m_{-1}$.
Note that Eq. (III.3) has the same form as Eq. (77) in Ref. [40]. Based on Eq. (III.3), the stability theorem in Ref. [40] can be extended as follows.
Theorem 1 The two-level explicit scheme for LCDE without the source term is stable ($|\lambda|\leq 1$) if and only if the following conditions holds,
@@
Remark 2 It is found from Eq. (71c) that $[(a_{k}-a_{-k})\bar{u}_{j}^{-}]^{2}\leq(a_{k}+a_{-k})\bar{u}_{j}^{+}$, even if $(a_{k}+a_{-k})\bar{u}_{j}^{+}=0$. If any $(a_{k}+a_{-k})\bar{u}_{j}^{+}=0$, then $(a_{k}-a_{-k})\bar{u}_{j}^{-}=0$, and thus Eq. (71c) follows by summing only over those $j$ and $k$ for which $(a_{k}+a_{-k})\bar{u}_{j}^{+}>0$.
@@
$\lambda=\frac{1+\theta(A+\textbf{i}B)}{1-(1-\theta)(A+\textbf{i}B)},$ (75)
@@
Theorem 2 The two-level $\theta$-scheme (Eq. (II.2)) for LCDE without the source term is (1) unconditional stable ($|\lambda|\leq 1$) for $20\leq\theta\leq 1/$ 2, (2) conditional stable for $\theta>1/2$ if Eq. (71a) holds.
Further, taking $n=2$ in Eq. (67) we can obtain the following characteristic polynomial of a general three-level scheme,
$p(\lambda)=\lambda^{2}-p_{1}\lambda-p_{0},p_{0}=\frac{A_{-2}}{1-A_{0}},p_{1}=\frac{A_{-1}}{1-A_{0}},$ (78)
where $A_{0},A_{-1}$, and $A_{-2}$ are defined by Eq. (65b). Then one can derive the following result from Schur-Cohn theorem [Miller1971].
Theorem 3 The three-level MMFD scheme is stable if and only if the following conditions holds,
$|p_{1}+\bar{p}_{1}p_{0}|\leq 1-|p_{0}|^{2},$ (79)
where $\bar{p}_{1}$ is the conjugate of $p_{1}$.
Some corollaries of Theorem 3 can be obtained for the specific three-level MMFD scheme. In the following, we take $A_{0}=0$ in Eq. (78), and two explicit schemes are considered.
For the TRT-LB model with $s^{+}+s^{-}=2$ as mentioned above, we have $p_{0}=s^{-}-1,~{}p_{1}=(2-s^{-})A-\textbf{i}s^{-}B$, where
$A=\sum_{j}\bar{u}_{j}^{+}\cos(\xi_{j}),B=\sum_{j}\bar{u}_{j}^{-}\sin(\xi_{j}),$ (80)
then Eq. (79) becomes $|A-\textbf{i}B|\leq 1$, and we obtain
Corollary 1 The TRT LB model with $s^{+}+s^{-}=2$ is stable if and only if
$A^{2}+B^{2}\leq 1.$ (81)
A sufficient condition of Eq. (81) was given in Ref. [22], which leads to the following corollary.
Corollary 2 The TRT LB model with $s^{+}+s^{-}=2$ is stable if the following conditions hold,
$(1)~{}\bar{u}_{0}^{+}\geq 0,\bar{u}_{j}^{+}>0,\forall j>0.~{}~{}(2)~{}\sum_{j\geq 1}\frac{(\bar{u}_{j}^{-})^{2}}{\bar{u}_{j}^{+}}\leq 1.$ (82)
Remark 3 Similarly to Remark 2 , if $\bar{u}_{j}^{+}=0$, then $\bar{u}_{j}^{-}=0$, and thus Eq. (82) follows by summing only over those $j$ for which $\bar{u}_{j}^{+}>0$.
For the three-level MMFD for WE mentioned above, we have $p_{0}=-1,~{}p_{1}=2-\bar{A}$, where
$\bar{A}=2\gamma\sum_{j}\bar{u}_{j}^{+}(1-\cos(k\xi_{j}))=4\gamma\sum_{j}\bar{u}_{j}^{+}\sin^{2}(k\xi_{j}/2),~{}\gamma=\frac{\alpha}{\beta c_{s}^{2}},$ (83)
then Eq. (79) is always true.
Now, to simplify the conditions in Eqs. (71) and (82), we consider the linear EDF. Taking $\mathbf{D}_{0}=\mathbf{I},\beta=1,\bar{\lambda}=0$ in Eq. (62), the linear EDF $f_{j}^{eq}$ with $\bar{u}_{j}^{+}=\omega_{j}>0$ and $\bar{u}_{j}^{-}=\omega_{j}\frac{\mathbf{c}_{j}\cdot\mathbf{u}}{c_{s}^{2}}$ is obtained, and
@@
The equivalent form of Eq. (60) can be further derived from Eq. (16) as
@@
For the two-level explicit scheme, Eq. (III.3) can be simplified by
@@
where Eq. (III.1), Eq. (42) and Eq. (85b) have been used. Note that Eq. (89b) is equivalent to Eq. (85b) or (86). When $m=1$, Eq. (89a) is simplified as
@@
and the stability condition in Eq. (85) becomes
@@
and the stability condition in Eq. (91) is equivalent to
@@
One can find that the condition in Eq. (96b) is much stronger than that in Eq. (97) for a large $d$.
@@ III.4 The derivation of the moment equations for NSEs using DTE method
where $\mathbf{M}_{k}$ and $\mathbf{M}_{kF}$ ($k\geq 0$) are the $k$-th moments of $f_{j},f_{j}^{eq}$ and $F_{j}$, respectively. Additionally, Eq. (98a) gives the following moments of nonequilibrium distribution function,
@@
Similarly, using DTE method we can derive the moment equations from Eq.(1) or Eq. (5). Before doing that, we first use the double-distribution based scheme (5) with the form of moments:
@@
in which two different set of parameters, $\{a_{kq},b_{kq}\}$ and $\{\bar{a}_{kq},\bar{b}_{kq}\}$, are used.
Using Taylor expansion to Eq.(100), we have
@@
Taking $\bar{A}_{00}=A_{00}=1$ in Eq. (101), we obtain a general moment equation for NSEs,
@@
Note that Eq. (103a) is same as Eq. (32). In what follows, we only discuss Eq.(103b) , and have the following equations for $N=1,2$,
@@
the moment equation at the leading order of $O(\Delta t^{2})$ can be derived from Eq. (106b),
@@
Furthermore, according to the NSEs, and choosing the moments and parameters in Eqs. (41) and (109) properly, we can correctly recover the NSEs from Eqs. (41) (or (110)) and (109). For example, taking
@@
in Eqs. (110) and (109) leads to the general NSEs,
@@
where the auxiliary source term $\mathbf{M}_{30}$ contains the additional term caused by the anisotropy of lattice tensor in a given lattice model, as mentioned in Ref. [17]. If an isotropic lattice is used, one can take $\mathbf{M}_{30}=\mathbf{0}$.
@@
therefore, the incompressible NSEs are recovered from Eqs. (113) and (112(b)) in the weak compressible limit, and $\mathbf{M}_{2F}$ can be simplified [17].